OpenStudy (anonymous):

Solve for the variable in the equation. 5x/7-1/14=9/28

OpenStudy (amistre64):

start celaring fracions by multiplying thru by the smallest one

OpenStudy (amistre64):

clearing that is

OpenStudy (amistre64):

that method is self simplifying

OpenStudy (anonymous):

So like starting with 1 over 14 first and multiply it by what?

OpenStudy (amistre64):

7 14 and 28; start by multiplying it all by 7, then simplify the rest and pick the next smallest denom

OpenStudy (anonymous):

Im not sure i quite understand.

OpenStudy (amistre64):

\[\frac{5x}{7}-\frac{1}{14}=\frac{9}{28}\] \[\frac{5x(7)}{7}-\frac{1(7)}{14}=\frac{9(7)}{28}\] \[5x-\frac{1}{2}=\frac{9}{4}\] \[(2)5x-\frac{1(2)}{2}=\frac{9(2)}{4}\] \[10x-1=\frac{9}{2}\] \[10(2)x-1(2)=\frac{9(2)}{2}\] \[20x-2=9\]

OpenStudy (anonymous):

So would that be 9/14?

OpenStudy (amistre64):

maybe, 11/20 tho might be better

OpenStudy (anonymous):

Oh haha i feel stupid lolz thanks your really smart, are you a student or an adult?

OpenStudy (amistre64):

im a student AND an adult .... getting my BA in math

OpenStudy (anonymous):

Wow good job;) I missed a month of school due to a car crash. And i havent been able to attend and now i dont get anything so i have a ton more questions that still need to be asked

OpenStudy (amistre64):

well, post them up on the left, one at a time so that they can be studied thru ... :) even if the intent is to get an answer, the studying part helps to solidify some of the concepts hopefully

OpenStudy (amistre64):

and if anyone ever just spouts out an answer, ask for proof cause they could be mistaken ;)

OpenStudy (anonymous):

Haha yeah, im glad you dont just tell me the answer to be perfectly honest with you, the people the other day and some of last night would just spout out answers to me and no one would talk me through it so i had no idea of what i was doing...but thanks:)

OpenStudy (anonymous):

Thanks for the medal:)